A torold ‏Has 750 turns the radius of the circle through the centre of its winding is 5 cm what current in the winding will produce in magnetic field of 1.8x10-3 tangent to this central circle


1
Expert's answer
2021-02-11T17:10:19-0500

"2\pi rB=\mu_0 NI,"

"I=\frac{2\pi rB}{\mu_0 N},"

"I=\frac{2\pi \cdot 0.05\cdot1.8\cdot 10^{-3}}{4\pi\cdot 10^{-7}\cdot 750}=0.6~\text{A}."